OpenStudy (anonymous):

Limit as x->1^- (x-1)/(absolute value(x^2-1)) How is that -1/2 instead of 1/2....

OpenStudy (agreene):

\[\lim_{x\rightarrow 1}(\frac{x-1}{|x^2-1|})\] is this the question?

OpenStudy (anonymous):

1^-

OpenStudy (agreene):

Ah, the issue is the jump/hole at x=1 from the left, at 1 y=-1/2 while, from the right at 1 y=1/2

OpenStudy (anonymous):

Ahhhhhhh its x-1/-(x^2-1) because x<1

OpenStudy (agreene):

indeed